# Weakly Supervised Object Segmentation by Background Conditional Divergence

### v1

@@ Abstract
As a computer vision task, automatic object segmentation remains challenging in specialized image domains without massive labeled data, such as synthetic aperture sonar images, remote sensing, biomedical imaging, etc. In any domain, obtaining pixel-wise segmentation masks is expensive. In this work, we propose a method for training a masking network to perform binary object segmentation using weak supervision in the form of image-wise presence or absence of an object of interest, which provides less information but may be obtained more quickly from manual or automatic labeling. A key step in our method is that the segmented objects can be placed into background-only images to create realistic, images of the objects with counterfactual backgrounds. To create a contrast between the original and counterfactual background images, we propose to first cluster the background-only images, and then during learning create counterfactual images that blend objects segmented from their original source backgrounds to backgrounds chosen from a targeted cluster. One term in the training loss is the divergence between these counterfactual images and the real object images with backgrounds of the target cluster. The other term is a supervised loss for background-only images. While an adversarial critic could provide the divergence, we use sample-based divergences. We conduct experiments on side-scan and synthetic aperture sonar in which our approach succeeds compared to previous unsupervised segmentation baselines that were only tested on natural images. Furthermore, to show generality we extend our experiments to natural images, obtaining reasonable performance with our method that avoids pretrained networks, generative networks, and adversarial critics. The basecode for this work can be found at https://github.com/bakerhassan/WSOS.
@@ 1 Introduction
Automatic foreground object segmentation is a computer vision task with a variety of applications in many fields such as medicine, autonomous driving, and defense (Ma et al., 2024; Song et al., 2024; Lee et al., 2024; Cao et al., 2024; Gao et al., 2024; Marinov et al., 2024; Kamtam et al., 2025). Supervised machine learning approaches are limited by the availability of supervised training annotations, such as object masks or bounding boxes, which in specialized domains are expensive to obtain (Russakovsky et al., 2015; Bakas et al., 2017; Sohn et al., 2020). Weak supervision in the form of image-level labels is often sufficient to guide segmentation. In particular, a well-trained classifier can provide class-activation maps to guide segmentation (Kumar Singh & Jae Lee, 2017; Wei et al., 2017; Zhang et al., 2018; Li et al., 2018; Xu et al., 2022; Xie et al., 2022). Prior work on unsupervised learning for object segmentation build off either pretrained classifiers, or the underlying feature extraction backbone networks (Remez et al., 2018; Li et al., 2023; Dosovitskiy et al., 2021; Hamilton et al., 2022; Zhang et al., 2018; Xie et al., 2022). However, for specialized imaging domains, such as sonar and medical images, there are not large crowd-sourced labeled datasets, and the limited datasets that are available differ in sensors and contexts that make it difficult to pretrain useful models. Thus, we consider the weakly supervised task for which no pretrained networks are available, where the dataset consists of two sets of images: one labeled as having an object (i.e., composite images) and the other as having no object (i.e., background-only images), which can be obtained through a combination of expert annotation, supervised image-level classification, and human-in-the-loop active learning.
@@
Our goal is to train a neural network to generate binary pixel-wise segmentation masks that separate salient foreground objects from the background. A motivating possibility for such a network is the ability to generate synthetic images composed of an object extracted from one image and background pixels from another image (Ostyakov et al., 2018). This synthetic image, created by alpha-blending a composite image and a background-only image, is referred to as a counterfactual image, as the background is counterfactual to the original background. While the automatic creation of counterfactual images is independently useful for augmenting training datasets to resolve spurious correlations between objects and their background, these counterfactual images are key to our approach as the network is trained to minimize the statistical divergence between the distributions of real images and the counterfactual images. [^1] We assume the object can be placed on an independently chosen background. This assumption is valid in underwater sonar where a man-made object’s location is largely independent of the background. However, in natural imagery there is much stronger dependence between the object and location within the image, due to depth and context, a problem addressed in other work (Zhu et al., 2023). To minimize this loss, the learned masks must generate counterfactual images that are statistically similar to the real images. However, without further specification this loss is minimized by a degenerate masking that segments the entire image as the object such that after alpha-blending the output is the original input images. Other work on unsupervised segmentation, which lacks the additional set of background-only images, has used the generators and discriminators underlying generative adversarial networks (GANs) along with additional loss terms to avoid such degenerate solutions (Bielski & Favaro, 2019; Sauer & Geiger, 2021; Zou et al., 2023; Yang et al., 2022b). Another solution proposed by Ostyakov et al. (2018) is to use a generative network to inpaint missing portions of the background after applying the complement mask, which should then be background only, and then calculate the divergence between inpainted images and real background-images.
@@
The differences between our approach and the majority of the other work reviewed in Section 2 can be summarized in the following:
@@
In particular, Table 4 shows that the related work suffer from at least one undesired attribute of requiring: pretrained models, adversarial training which is known for its instability, labeled data (e.g., class information), or extensive hyperparameter search for a complicated loss function combination (i.e., not minimalist).
@@
    Toy datasets created by combining canonical datasets (e.g., MNIST, Fashion MNIST (Xiao et al., 2017), and dSprites (Matthey et al., 2017)) with textured backgrounds (Brodatz, 1966). We show that existing unsupervised segmentation benchmarks (Bielski & Favaro, 2019; Xie et al., 2022) fail to converge to an acceptable segmentation on these datasets.
@@
    Side-Scan Sonar (SSS) images containing various foreground objects such as shipwrecks and rocks (Sethuraman et al., 2024). We show that our methodology provides finer-grained details for foreground objects compared to the available ground truth.
@@
    Synthetic Aperture Sonar (SAS) images (Cobb & Zare, 2014), where artificial objects are pasted onto real sonar backgrounds.
@@
    Natural images from the CUB-200-2011 dataset (Wah et al., 2011), which consists of bird images without background-only samples. Our approach performs reasonably well on this dataset, achieving an average precision of up to 50%.
@@ 2 Related Work
Our weakly supervised foreground object segmentation is most closely related to prior work on unsupervised foreground object segmentation (Bielski & Favaro, 2019; Sauer & Geiger, 2021; Zou et al., 2023; Yang et al., 2022b). These methods either depend on adversarial training (Ostyakov et al., 2018; Remez et al., 2018; Bielski & Favaro, 2019; Chen & Konukoglu, 2018; Singh et al., 2019; He et al., 2022; Yang et al., 2022a; Zou et al., 2023; Sauer & Geiger, 2021; Li et al., 2023), make assumptions about the foreground object size (Bielski & Favaro, 2019; He et al., 2022; Yang et al., 2022b; Zou et al., 2023; Sauer & Geiger, 2021), depend on a given initial mask (Li et al., 2023), need an extensive hyperparameters search (Li et al., 2018; Remez et al., 2018; Bielski & Favaro, 2019; Singh et al., 2019; He et al., 2022; Yang et al., 2022b; Zou et al., 2023; Sauer & Geiger, 2021; Hung et al., 2019; Savarese et al., 2021), depend on pre-trained models or foreground labels (e.g., birds classes) (Remez et al., 2018; Sauer & Geiger, 2021; Li et al., 2023; Dosovitskiy et al., 2021; Hamilton et al., 2022; Wang et al., 2023b; Zhang et al., 2018; Li et al., 2018; Hou et al., 2018; Jiang et al., 2019; Chang et al., 2020; Lee et al., 2021; Wu et al., 2021; Xie et al., 2022; Hung et al., 2019), require stochastic generation or complicated combinations of losses to avoid degenerate solutions (Ostyakov et al., 2018; Li et al., 2018; Remez et al., 2018; Bielski & Favaro, 2019; Hung et al., 2019; Zou et al., 2023; Yang et al., 2022b). Our proposed method, on the other hand, does not depend on any of that. The difference is that we assume access to background-only images that follow the same distribution of backgrounds in the composite images. (Ostyakov et al. (2018) also use this assumption but have adversarial discriminator along with an enhancement and inpainting network.) While purely unsupervised learning without any background-only images is even more challenging, in practice background-only images can be obtained from the surrounding of images with an object, and are naturally obtained through the weak labeling process.
@@
Furthermore, the majority of related work has focused on natural images, where there is often visible differences between objects and background due to the optics. Additionally, in other imaging domains, there may also be a lack of pretrained models. In the case of underwater sonar imagery, which is a specialized and limited domain, even more so than standardized medical imaging (Ma et al., 2024), with datasets limited by the diversity of objects and sensors (Valdenegro-Toro et al., 2021; Preciado-Grijalva et al., 2022). Nonetheless, self-supervised pretraining before classification is possible (Preciado-Grijalva et al., 2022). We found that the majority of the work on unsupervised segmentation of sonar imagery is concerned with sea floor segmentation (Sun et al., 2022; Abu & Diamant, 2017; Yao et al., 2000) or object detection (Abu & Diamant, 2018; Wei et al., 2024). One recent work did investigate generating new sonar imagery using a mask-conditioned diffusion model (i.e., counterfactual images); however they did not explicitly address segmentation (Yang et al., 2024).
@@
A description of related work is included in Appendix A, and summarized in Table 4, which notes which works depend on pretrained models, GAN training, or data labels (e.g., foreground labels). The table highlights that prior work has at least one of these undesirable criteria, which makes it hard to be applied to data-scarce domains such as sonar images.
@@ 3 Methodology
Our methodology trains a masking network by using its output to alpha-blend the input composite image with an independent background-only image creating a counterfactual image. The training loss consists of two terms; the first is a measure of the statistical divergence between counterfactual and real composite images, and the second is a pixel-wise supervised loss that encourages the alpha-blended output to match the input for background-only images. To avoid degenerate solutions we rely on conditioning the divergence based on different background clusters, which we derive from an auto-encoder or contrastive-learning based representation. Additionally, we consider an extra loss for the masks of composite images that forces a portion of the mask to not be background, which is applicable if the object size is known a priori. Our approach requires minimal hyper-parameter tuning, namely the number of clusters, epochs, and batch size. For architectures we use a U-Net (Ronneberger et al., 2015b), which is a standard neural network architecture, as the masking network and minimal auto-encoder or encoder-only architectures for the latent representation of backgrounds. Figure 1 summarizes the approach.
@@ 3.1 Problem Formulation
where $R\in\mathcal{R}$ is a background-only image, $F\in\mathcal{I}$ is an image of an object paired with a binary mask $M\in\{0,1\}^{m\times n}$ ($M_{ij}=1$ if the pixel is object and $M_{ij}=0$ if the pixel is background), and $\mathbf{1}$ is a $m\times n\times c$-array of ones. The unsupervised segmentation goal is to learn to infer $M$ given only $X_{R}$.
@@
Probabilistically, as a random variable $X_{R}\sim\mathbb{P}_{X}$ is a function of the dependent random variables $(F,M,R)\sim\mathbb{P}_{M,F,R}$, since the object and the background are dependent. However, we assume $R\sim\mathbb{P}_{R}$ is independent of $(F,M)\sim\mathbb{P}_{M,F}$. Under this assumption, one can generate
$X_{R^{\prime}}=M\circ X_{R}+(\mathbf{1}-M)\circ R^{\prime},$ (2)
which is an image with the same object (and mask), but a different background $R^{\prime}\sim\mathbb{P}_{R}$. While $X_{R}$ and $R^{\prime}$ represent real images, $X_{R^{\prime}}$ is a counterfactual that is a function of $M$. If the independence assumptions hold, $X_{R^{\prime}}\sim\mathbb{P}_{X}$.
@@
A masking network is a function $M_{\theta}:\mathcal{I}\rightarrow[0,1]^{m\times n}$, parameterized by $\theta$, that models the relationship between $X_{R}$ and $M$. Unlike, supervised segmentation the latter is never observed. An ideal masking network would yield the true mask $M_{\theta^{*}}(X_{R})=M$ when applied to a composite image, and would yield an all-zeros mask $\mathbf{0}$ when applied to a background only image $M_{\theta^{*}}(R)=\mathbf{0}$. The output mask is used to alpha-blend a composite image with an independent background-only image $R^{\prime}$,
$\tilde{X}_{R^{\prime}}=M_{\theta}(X_{R})\circ X_{R}+(\mathbf{1}-M_{\theta}(X_{R}))\circ R^{\prime}\in\tilde{\mathcal{X}}\subset\mathcal{I}$ (3)
@@
The goal is then to make the distribution of $\tilde{X}_{R^{\prime}}$, $\mathbb{P}_{\tilde{X}}$ statistically indistinguishable from $\mathbb{P}_{X}$. However, directly minimizing a statistical divergence between $\mathbb{P}_{\tilde{X}}$ and $\mathbb{P}_{X}$ leads to a trivial solution where the masking network always produces a mask of all ones (all foreground),
$M_{\text{degen}}:X\mapsto\mathbf{1}=\operatorname*{arg\,min}_{M_{\theta}}D(\mathbb{P}_{\tilde{X}},\mathbb{P}_{X}),$ (4)
@@
since $\tilde{X}_{R^{\prime}}=M_{\text{degen}}(X_{R})\circ X_{R}+(\mathbf{1}-M_{\text{degen}}(X_{R}))\circ R^{\prime}=X_{R}\sim\mathbb{P}_{X}$, a problem also encountered and addressed in different ways in GAN-based approaches for unsupervised segmentation (Bielski & Favaro, 2019; Yang et al., 2022b; Zou et al., 2023) and weakly supervised segmentation (Ostyakov et al., 2018).
@@
To overcome this without the use of GANs, we create a distributional shift between $R$ and $R^{\prime}$ based on the background portion of the composite image $X_{R}$ and the background-only image $R^{\prime}$. The intuition is that restricting the counterfactuals to be generated from source backgrounds that are different than the backgrounds of the target composite images background, discourages masks from retaining any dissimilar background as that would increase the divergence. Specifically, we assume $(R,C_{b})\sim\mathbb{P}_{R,C_{b}}$ is a joint distribution over background images and a latent background class $C_{b}\in\{1,\dots,K\}$. The latent variable $C_{b}$ carries information about the background. Let $\mathbb{P}_{X|C_{b}=c}$ represent the distribution of composite images generated under the condition that the background image belongs to class $c\in\{1,\ldots,K\}$. When computing the divergence the goal is to target composite images with a specific background class $\mathbb{P}_{X|C_{b}=c}$, generating counterfactuals from independent background-only images of the same class $R_{c}\sim\mathbb{P}_{R|C_{b}=c}$ and sourcing composite images that do not have this class $X_{\neg c}=X_{R|C_{b}\neq c}\sim\mathbb{P}_{X|C_{b}\neq c}$, where $\mathbb{P}_{R|C_{b}\neq c}$ is the distribution over other background classes, such that $X_{c}\sim\mathbb{P}_{X|C_{b}=c}$ and $X_{\neg c}$ are independent random variables with different distributions. The masking network’s counterfactual image with background from class $c$ is
@@
The goal is to optimize the parameters $\theta$ of the masking network $M_{\theta}$ in terms of a the background conditional divergence between the distributions of $\mathbb{P}_{X|C_{b}=c}$ and $\mathbb{P}_{\tilde{X}_{c,\theta}}$,
@@
Due to the distinct distribution of $X_{\neg c}$ and $X_{c}$, the trivial masking solution will not minimize the divergence between $\mathbb{P}_{\tilde{X}_{c,\theta}}$ and $\mathbb{P}_{X|C_{b}=c}$, $M_{\text{degen}}(X_{\neg c})\circ X_{\neg c}+(\mathbf{1}-M_{\text{degen}}(X_{\neg c}))\circ R_{c}=X_{\neg c}$ is clearly not distributed according to $\mathbb{P}_{X|C_{b}=c}$.
Since we assume a sample of background-only images, we also train the masking network to determine if there is an object in the input image (by outputting an all zero mask). While a binary cross-entropy loss could be used per pixel, we use the squared error loss
[…]

### v2

@@ Abstract
As a computer vision task, automatic object segmentation remains challenging in specialized image domains without massive labeled data, such as synthetic aperture sonar images, remote sensing, biomedical imaging, etc. In any domain, obtaining pixel-wise segmentation masks is expensive. In this work, we propose a method for training a masking network to perform binary object segmentation using weak supervision in the form of image-wise presence or absence of an object of interest, which provides less information but may be obtained more quickly from manual or automatic labeling. A key step in our method is that the segmented objects can be placed into background-only images to create realistic images of the objects with counterfactual backgrounds. To create a contrast between the original and counterfactual background images, we propose to first cluster the background-only images and then, during learning, create counterfactual images that blend objects segmented from their original source backgrounds to backgrounds chosen from a targeted cluster. One term in the training loss is the divergence between these counterfactual images and the real object images with backgrounds of the target cluster. The other term is a supervised loss for background-only images. While an adversarial critic could provide the divergence, we use sample-based divergences. We conduct experiments on side-scan and synthetic aperture sonar in which our approach succeeds compared to previous unsupervised segmentation baselines that were only tested on natural images. Furthermore, to show generality we extend our experiments to natural images, obtaining reasonable performance with our method that avoids pretrained networks, generative networks, and adversarial critics. The code for this work can be found at https://github.com/bakerhassan/WSOS.
@@ 1 Introduction
Automatic foreground object segmentation is a computer vision task with a variety of applications in many fields such as medicine, autonomous driving, and defense (Ma et al., 2024; Song et al., 2024; Lee et al., 2024; Cao et al., 2024; Gao et al., 2024; Marinov et al., 2024; Kamtam et al., 2025). Supervised machine learning approaches are limited by the availability of supervised training annotations, such as object masks or bounding boxes, which in specialized domains are expensive to obtain (Russakovsky et al., 2015; Bakas et al., 2017; Sohn et al., 2020). Weak supervision in the form of image-level labels is often sufficient to guide segmentation. In particular, a well-trained classifier can provide class-activation maps to guide segmentation (Kumar Singh & Jae Lee, 2017; Wei et al., 2017; Zhang et al., 2018; Li et al., 2018; Xu et al., 2022; Xie et al., 2022). Prior work on unsupervised learning for object segmentation builds off either pretrained classifiers or the underlying feature extraction backbone networks (Remez et al., 2018; Li et al., 2023; Dosovitskiy et al., 2021; Hamilton et al., 2022; Zhang et al., 2018; Xie et al., 2022). However, for specialized imaging domains, such as sonar and medical images, there are not large crowd-sourced labeled datasets, and the limited datasets that are available differ in sensors and contexts that make it difficult to pretrain useful models. Thus, we consider the weakly supervised task for which no pretrained networks are available, where the dataset consists of two sets of images: one labeled as having an object (i.e., composite images) and the other as having no object (i.e., background-only images), which can be obtained through a combination of expert annotation, supervised image-level classification, and human-in-the-loop active learning.
@@
Our goal is to train a neural network to generate binary pixel-wise segmentation masks that separate salient foreground objects from the background. A motivating possibility for such a network is the ability to generate synthetic images composed of an object extracted from one image and background pixels from another image (Ostyakov et al., 2018). This synthetic image, created by alpha-blending a composite image and a background-only image, is referred to as a counterfactual image, as the background is counterfactual to the original background. While the automatic creation of counterfactual images is independently useful for augmenting training datasets to resolve spurious correlations between objects and their background, these counterfactual images are key to our approach as the network is trained to minimize the statistical divergence between the distributions of real images and the counterfactual images. [^1] We assume the object can be placed on an independently chosen background. This assumption is valid in underwater sonar where a man-made object’s location is largely independent of the background. However, in natural imagery there is much stronger dependence between the object and location within the image, due to depth and context, making it imperative to choose an appropriate background image and/or location within the background, a problem addressed in other work (Dvornik et al., 2018; Zhu et al., 2023). To minimize this loss, the learned masks must generate counterfactual images that are statistically similar to the real images. However, without further specification this loss is minimized by a degenerate masking that segments the entire image as the object such that after alpha-blending the output is the original input images. Other work on unsupervised segmentation, which lacks the additional set of background-only images, has used the generators and discriminators underlying generative adversarial networks (GANs) along with additional loss terms to avoid such degenerate solutions (Bielski & Favaro, 2019; Sauer & Geiger, 2021; Zou et al., 2023; Yang et al., 2022b). Another solution proposed by Ostyakov et al. (2018) is to use a generative network to inpaint missing portions of the background after applying the complement mask, which should then be background only, and then calculate the divergence between inpainted images and real background-images.
@@
The differences between our approach and the majority of the other work reviewed in Section 2 are the following:
@@
In particular, Table 6 shows that the related work suffers from requiring at least one of the following undesired attributes: pretrained models, adversarial training (which is known for its instability), labeled data (e.g., class information), or extensive hyperparameter search for a complicated loss function combination (i.e., not minimalist).
@@
    Toy datasets created by combining canonical object datasets (e.g., MNIST, Fashion MNIST (Xiao et al., 2017), and dSprites (Matthey et al., 2017)) with textures as backgrounds (Brodatz, 1966). We show that existing unsupervised segmentation benchmarks (Bielski & Favaro, 2019; Xie et al., 2022) fail to converge to an acceptable segmentation on these grayscale image datasets.
@@
    AI4Shipwrecks consists of side-scan sonar images of shipwrecks with other foreground objects such as large rocks (Sethuraman et al., 2024). We show that our methodology provides finer-grained details for foreground objects compared to the available ground truth.
@@
    SAS-Clutter consists of small synthetic aperture sonar (SAS) image snippets containing seafloor clutter objects such as crab traps, barrels, tires, etc. Results demonstrate meaningful segmentation. The dataset is not publicly available.
@@
    SAS+dSprites consists of real SAS images of various seafloor types (Cobb & Zare, 2014) with artificial shapes from dSprites (Matthey et al., 2017) as synthetic objects.
  5. 5.
    CUB-200-2011 consists of natural images of 200 species of birds (Wah et al., 2011).
@@ 2 Related Work
Our weakly supervised foreground object segmentation is most closely related to prior work on unsupervised foreground object segmentation (Bielski & Favaro, 2019; Sauer & Geiger, 2021; Zou et al., 2023; Yang et al., 2022b). These methods either depend on adversarial training (Ostyakov et al., 2018; Remez et al., 2018; Bielski & Favaro, 2019; Chen & Konukoglu, 2018; Singh et al., 2019; He et al., 2022; Yang et al., 2022a; Zou et al., 2023; Sauer & Geiger, 2021; Li et al., 2023), make assumptions about the foreground object size (Bielski & Favaro, 2019; He et al., 2022; Yang et al., 2022b; Zou et al., 2023; Sauer & Geiger, 2021), depend on a given initial mask (Li et al., 2023), need an extensive hyperparameters search (Li et al., 2018; Remez et al., 2018; Bielski & Favaro, 2019; Singh et al., 2019; He et al., 2022; Yang et al., 2022b; Zou et al., 2023; Sauer & Geiger, 2021; Hung et al., 2019; Savarese et al., 2021), depend on pretrained models or foreground labels (e.g., birds classes) (Remez et al., 2018; Sauer & Geiger, 2021; Li et al., 2023; Dosovitskiy et al., 2021; Hamilton et al., 2022; Wang et al., 2023b; Zhang et al., 2018; Li et al., 2018; Hou et al., 2018; Jiang et al., 2019; Chang et al., 2020; Lee et al., 2021; Wu et al., 2021; Xie et al., 2022; Hung et al., 2019), or require stochastic generation or complicated combinations of losses to avoid degenerate solutions (Ostyakov et al., 2018; Li et al., 2018; Remez et al., 2018; Bielski & Favaro, 2019; Hung et al., 2019; Zou et al., 2023; Yang et al., 2022b). Our proposed method, on the other hand, does not depend on any of these undesirable attributes. The difference between these prior methods and our work is that we assume access to background-only images that follow the same distribution of backgrounds in the composite images. (Ostyakov et al. (2018) also use this assumption but have an adversarial discriminator along with an enhancement and inpainting network.) While purely unsupervised learning without any background-only images is even more challenging, in practice background-only images can be obtained from the surrounding of images with an object and are naturally obtained through the weak labeling process.
@@
Furthermore, the majority of related work has focused on natural images where there is often visible differences between objects and background due to the optics. Additionally, in other imaging domains, there may also be a lack of pretrained models. In the case of underwater sonar imagery, which is a specialized and limited domain—even more so than standardized medical imaging (Ma et al., 2024)—datasets are limited by the diversity of objects and sensors (Valdenegro-Toro et al., 2021; Preciado-Grijalva et al., 2022). Nonetheless, self-supervised pretraining before classification is possible (Preciado-Grijalva et al., 2022). We found that the majority of the work on unsupervised segmentation of sonar imagery is concerned with sea floor segmentation (Sun et al., 2022; Abu & Diamant, 2017; Yao et al., 2000) or object detection (Abu & Diamant, 2018; Wei et al., 2024). One recent work did investigate generating new sonar imagery using a mask-conditioned diffusion model (i.e., counterfactual images); however, they did not explicitly address segmentation (Yang et al., 2024).
@@
A description of related work is included in Appendix A and summarized in Table 6, which notes which works depend on pretrained models, GAN training, or data labels (e.g., foreground labels). The table highlights that all prior work has at least one of these undesirable criteria, which makes it hard to apply any of these methods to data-scarce domains such as sonar images.
@@ 3 Methodology
Our methodology trains a masking network by using its output to alpha-blend the input composite image with an independent background-only image creating a counterfactual image. The training loss consists of two terms; the first is a measure of the statistical divergence between counterfactual and real composite images, and the second is a pixel-wise supervised loss that encourages the alpha-blended output to match the input for background-only images. To avoid degenerate solutions we rely on conditioning the divergence based on different background clusters, which we derive from an auto-encoder or contrastive-learning based representation. Additionally, we consider an extra loss for the masks of composite images that forces a portion of the mask to not be background, which is applicable if the object size is known a priori. Our approach requires minimal hyperparameter tuning, namely the number of clusters, epochs, and batch size. For architectures we use a U-Net (Ronneberger et al., 2015b), which is a standard neural network architecture, as the masking network and minimal auto-encoder or encoder-only architectures for the latent representation of backgrounds. Figure 1 summarizes the approach, and Table 1 details the variables in the formulation.
@@ 3.1 Problem Formulation
Table 1: Summary of random variable notation used in the problem formulation.
Notation Description
$(F,M,R)\sim\mathbb{P}_{M,F,R}$ An object image $F\in\mathcal{I}$, a binary mask $M\in\{0,1\}^{m\times n}$, and a background image $R\in\mathcal{R}\subset\mathcal{I}$ jointly sampled.
$R\sim\mathbb{P}_{R}$ A background image sampled from the background distribution.
$X_{R}\sim\mathbb{P}_{X}$ A composite image constructed from $F$, $M$, and background $R$.
$X_{R^{\prime}}$ An ideal counterfactual composite image constructed from $F$ and $M$ (or equivalently $X_{R}$ and $M$) but with background $R^{\prime}\sim\mathbb{P}_{R}$. If object-background independence holds $X_{R^{\prime}}\sim\mathbb{P}_{X}$, then $X_{R^{\prime}}\stackrel{{\scriptstyle d}}{{=}}X_{R}$, where $\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution.
$M_{\theta}$ Masking network $M_{\theta}:\mathcal{I}\rightarrow[0,1]^{m\times n}$.
$\tilde{X}_{R^{\prime}}\sim\mathbb{P}_{\tilde{X}_{\theta}}$ A counterfactual image produced from $X_{R}$ and $R^{\prime}$ using the mask $M_{\theta}(X_{R})$. Ideally, $\tilde{X}_{R^{\prime}}\stackrel{{\scriptstyle d}}{{=}}X_{R^{\prime}}\stackrel{{\scriptstyle d}}{{=}}X_{R}$, assuming object-background independence.
$M_{\text{degen}}$ A degenerate masking function that maps any composite image to a mask of all ones, i.e., $M_{\text{degen}}(X)=\mathbf{1}\in\{1\}^{m\times n},\quad\forall X\in\mathcal{X}$.
$C_{b}\sim\text{Cat}(1,\dots,K)$ A random variable for the background cluster label, drawn from a categorical distribution over $K$ clusters.
$X_{c}\sim\mathbb{P}_{X|C_{b}=c}$ A composite image with a background from cluster $c$.
$R_{c}\sim\mathbb{P}_{R|C_{b}=c}$ A background image from cluster $c$.
$X_{\neg c}\sim\mathbb{P}_{X|C_{b}\neq c}$ A composite image with a background not from cluster $c$.
$\tilde{X}_{c}\sim\mathbb{P}_{\tilde{X}_{c,\theta}}$ Counterfactual image produced from $X_{\neg c}$ and $R_{c}$ using the mask $M_{\theta}(X_{\neg c})$. Ideally, $\tilde{X}_{c}\stackrel{{\scriptstyle d}}{{=}}X_{c}$.
@@
where $\circ$ denotes an elementwise product with broadcasting across channels, $R\in\mathcal{R}$ is a background-only image, $F\in\mathcal{I}$ is an image of an object paired with a binary mask $M\in\{0,1\}^{m\times n}$ ($M_{ij}=1$ if the pixel corresponds to the object and $M_{ij}=0$ if the pixel corresponds to background), and $\mathbf{1}$ is a $m\times n$-array of ones. The unsupervised segmentation goal is to learn to infer $M$ given only $X_{R}c\sim\mathbb{P}_{X}$, where $\mathbb{P}_{X}$ is the distribution of composite images.
@@
Probabilistically, as a random variable $X_{R}\sim\mathbb{P}_{X}$ is a function of the dependent random variables $(F,M,R)\sim\mathbb{P}_{M,F,R}$, where the object and mask are dependent, and the object and the background may be dependent. However, we assume $R\sim\mathbb{P}_{R}$ is independent of $(F,M)\sim\mathbb{P}_{M,F}$. Under this assumption, one can generate a new composite image $X_{R^{\prime}}$ containing the same object $M\circ X_{R}=M\circ F$ but with a different independent background $R^{\prime}\sim\mathbb{P}_{R}$:
$X_{R^{\prime}}=M\circ X_{R}+(\mathbf{1}-M)\circ R^{\prime}.$ (2)
While $X_{R}$ and $R^{\prime}$ represent real images, $X_{R^{\prime}}$ is an ideal counterfactual. If the independence assumptions hold, $X_{R^{\prime}}\sim\mathbb{P}_{X}$.  [^2] In domains with strong dependence between objects and background, in order to create the counterfactual $X_{R^{\prime}}$, the background $R^{\prime}$ would need to be conditional on the object $F,M$, a problem considered in other works (Fang et al., 2019; Dvornik et al., 2018; Zhu et al., 2023).
@@
A masking network is a function $M_{\theta}:\mathcal{I}\rightarrow[0,1]^{m\times n}$, parameterized by $\theta$, that models the relationship between $X_{R}$ and MM. [^3] While an ideal masking network would have binary output, to facilitate training the unit interval is allowed. Unlike supervised segmentation, the latter is never observed. An ideal masking network would yield the true mask $M_{\theta^{*}}(X_{R})=M$ when applied to a composite image and would yield an all-zeros mask $\mathbf{0}$ when applied to a background only image $M_{\theta^{*}}(R)=\mathbf{0}$. The output mask is used to alpha-blend a composite image $X_{R}\sim\mathbb{P}_{X}$ with an independent background-only image $R^{\prime}\sim\mathbb{P}_{R}$,
$\tilde{X}_{R^{\prime}}=M_{\theta}(X_{R})\circ X_{R}+(\mathbf{1}-M_{\theta}(X_{R}))\circ R^{\prime}\sim\mathbb{P}_{\tilde{X}_{\theta}}.$ (3)
@@
The goal is then to make the counterfactual distribution $\mathbb{P}_{\tilde{X}_{\theta}}$ statistically indistinguishable from the composite image distribution $\mathbb{P}_{X}$. However, directly minimizing a statistical divergence between $\mathbb{P}_{\tilde{X}}$ and $\mathbb{P}_{X}$ leads to a trivial solution $M_{\text{degen}}:X\mapsto\mathbf{1},\quad\forall X\in\mathcal{X}$, where the masking network always produces a mask of all ones (indicating all foreground) so that each generated counterfactual image is the same as its input composite image. That is,
$M_{\text{degen}}=\operatorname*{arg\,min}_{M_{\theta}}D(\mathbb{P}_{\tilde{X}_{\theta}},\mathbb{P}_{X}),$ (4)
@@
based on the fact that $\tilde{X}_{R^{\prime}}=M_{\text{degen}}(X_{R})\circ X_{R}+(\mathbf{1}-M_{\text{degen}}(X_{R}))\circ R^{\prime}=X_{R}\sim\mathbb{P}_{X}$, a problem also encountered and addressed in different ways in GAN-based approaches for unsupervised segmentation (Bielski & Favaro, 2019; Yang et al., 2022b; Zou et al., 2023) and weakly supervised segmentation (Ostyakov et al., 2018).
@@
To overcome this, we create a distributional shift between the random backgrounds $R$ found in the composite images and the random background-only images. We restrict the counterfactuals $\tilde{X}_{R^{\prime}}$ to be generated from backgrounds $R^{\prime}$ that are different from the backgrounds $R$ of the composite images $X_{R}$. The divergence to be minimized is then computed only between distributions of the same background. This discourages masks from retaining the original (composite) backgrounds $R$, as this would increase the (background-conditional) divergence. Specifically, we assume that all backgrounds belong to a latent background class $C_{b}\in\{1,\dots,K\}$, and that there exists a joint distribution $(R,C_{b})\sim\mathbb{P}_{R,C_{b}}$ over background images and latent background class. Let $\mathbb{P}_{X|C_{b}=c}$ represent the distribution of composite images generated under the condition that the background image belongs to class $c\in\{1,\ldots,K\}$. To target composite images with a specific background class $\mathbb{P}_{X|C_{b}=c}$, the goal is to generate counterfactuals from independent background-only images of the same class $R_{c}\sim\mathbb{P}_{R|C_{b}=c}$ and composite images that do not have this class $X_{\neg c}=X_{R|C_{b}\neq c}\sim\mathbb{P}_{X|C_{b}\neq c}$, where $\mathbb{P}_{R|C_{b}\neq c}$ is the distribution over other background classes, such that $X_{c}\sim\mathbb{P}_{X|C_{b}=c}$ and $X_{\neg c}$ are independent random variables with different distributions. The masking network’s counterfactual image with background from class $c$ is
@@
We optimize the parameters $\theta$ of the masking network $M_{\theta}$ in terms of a background conditional divergence between the distributions of $\mathbb{P}_{X|C_{b}=c}$ and $\mathbb{P}_{\tilde{X}_{c,\theta}}$,
@@
Due to the distinct distribution of $X_{\neg c}$ and $X_{c}$, the trivial masking solution will not minimize the divergence between $\mathbb{P}_{\tilde{X}_{c,\theta}}$ and $\mathbb{P}_{X|C_{b}=c}$, $M_{\text{degen}}(X_{\neg c})\circ X_{\neg c}+(\mathbf{1}-M_{\text{degen}}(X_{\neg c}))\circ R_{c}=X_{\neg c}$ is clearly not distributed according to $\mathbb{P}_{X|C_{b}=c}$. We summarize the notation in Table 1.
Since we assume a sample of background-only images, we also train the masking network to determine if there is no object in the input image (by outputting an all-zero mask). While a per-pixel binary cross-entropy loss could be used, we use the squared error loss
$\mathcal{L}_{\text{Bg}}(\theta)=\mathbb{E}_{R}\Big[\sum_{ij}[M_{\theta}(R)]_{ij}^{2}\Big]=\mathbb{E}_{R}\left[\lVert M_{\theta}(R)\rVert_{F}^{2}\right].$ (7)